OpenStudy (anonymous):

determine convergence or divergence. for converging series, state absolute or conditional (explain why): sum(k=2 -> infinity) [(-1)^k/(k^2-1)]

OpenStudy (zarkon):

you can actually find the sum

OpenStudy (zarkon):

though a limit comparison test will tell you quickly it converges

OpenStudy (anonymous):

Use alternating series test to say: \[a_{k+1}<a_k \] For sufficiently large n. \[a_k \ge 0\] For all n. \[\lim_{k \rightarrow \infty} a_k \rightarrow 0\] Therefore it converges by the alternating series test. So it is conditionally convergent. Now to find it it is absolutely convergent consider the summation: \[\sum_{k=2}^{\infty} \frac{1}{k^2-1}\] By limit comparison with \[\sum_{k=2}^{\infty} \frac{1}{k^2}\] This series converges so we can say that the original series converges based on the limit comparison test.

OpenStudy (anonymous):

Sufficiently large k. For all k. Not "n".

OpenStudy (anonymous):

thanks.. can you help me out with finding if it converges absolutely or conditionaly

OpenStudy (anonymous):

Well, conditionally means it converges ONLY with the (-1)^k, and absolutely means it converges regardless. So since it converges in both cases it converges absolutely.

OpenStudy (anonymous):

@ Zarkon

OpenStudy (anonymous):

@Zarkon

OpenStudy (zarkon):

yes?

OpenStudy (anonymous):

can you help me out with finding if it converges absolutely or conditionaly

OpenStudy (zarkon):

it is absolutely convergent by the limit comparison test

OpenStudy (anonymous):

is it because the absolute value of zero is zero?

OpenStudy (zarkon):

no

OpenStudy (anonymous):

oops, is it because the absolute value of (-1)^k/(k^2-1) is also zero

OpenStudy (zarkon):

\[\lim_{k\to\infty}\frac{k^2}{k^2-1}=1\] the limit is positive finite so \[\sum_{k=2}^{\infty}\frac{1}{k^2-1}\] converges since \[\sum_{k=2}^{\infty}\frac{1}{k^2}\] converges

OpenStudy (zarkon):

malevolence19 provided you with what you needed...though there is no need for the alternating series test (in this case)

OpenStudy (anonymous):

thanks..